OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches negative four of quantity x squared minus sixteen divided by quantity x plus four

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Directrix please help

OpenStudy (anonymous):

(x^2-16) / (x+4) = (x+4)(x-4) / (x+4)

OpenStudy (anonymous):

you are just factoring so far so we can use -4 as the x value to solve for the limit

OpenStudy (anonymous):

if insert -4 into the equation before factoring we get a denominator of 0 which we cant use

OpenStudy (anonymous):

so factor the top and cancel ie... for (x+4)(x-4) /(x+4) you can cancel the (x+4)'s can you are left with (x-4)

OpenStudy (anonymous):

inserting the limit of -4 into the equation you have f(-4) = (-4-4) =-8

OpenStudy (anonymous):

so the limit of f(x) as x approaches -4 = -8